# Generative Artificial Intelligence for Beamforming in Low-Altitude Economy

### v1

@@ III-C1 Communication Tasks
  • Channel Modeling and Estimation: Compared to conventional communication scenarios, LAE places higher demands on channel modeling and estimation due to its complex and highly dynamic environment. Specifically, factors such as adverse weather conditions, obstruction by buildings or terrain features, interference from surrounding aerial and ground-based nodes, and the rapid mobility of LAPs significantly impact signal quality. These external factors often lead to severe signal attenuation and pronounced multipath propagation effects, thus posing substantial challenges to accurate channel modeling and estimation. Although conventional beamforming techniques can adapt to certain channel variations, they fall short in effectively addressing the rapid temporal and spatial fluctuations introduced by these external environmental factors. Therefore, there is a clear need for innovative approaches, and GenAI offers promising solutions by leveraging advanced data-driven frameworks to improve channel modeling and estimation.
    For instance, Zhang et al. [8] developed a GAN-based model to improve channel estimation in millimeter-wave communications for low-altitude unmanned aerial vehicle (UAV) networks. Specifically, the model allows each UAV to independently train channel models for specific beamforming directions, with samples shared through a collaborative framework to build a generalized model, which improves estimation accuracy. Additionally, Rasheed et al. [9] proposed a distributed GAN model based on long short-term networks for low-altitude maritime UAV communications. Specifically, UAVs utilize the model to extract real-time CSI and generate synthetic samples, which improve the accuracy of channel state estimation.
@@ III-C2 Performance Metrics
  • Energy Consumption: In LAE senarios, the frequent beam redirection by LAPs, interference suppression between high-density nodes, and multi-network handovers require continuous power consumption, thus leading to a significant increase in energy usage compared to conventional communication environments. However, conventional beamforming algorithms rely on pre-set power distribution strategies, which hinder the ability to balance communication quality and energy efficiency in dynamic topology and channel fluctuation scenarios. GenAI can learn the relationship between environmental features and energy consumption online, autonomously generating beam control strategies that optimize energy efficiency and overcome the limitations of traditional static approaches.
@@ III-D Lessons Learned
    GenAI models fundamentally enhance beamforming adaptability in dynamic environments through real-time parameter generation and high-fidelity channel modeling. This capability directly addresses the rapid spatial-temporal variations and interference challenges specific to low-altitude scenarios while maintaining robust performance under mobility constraints and environmental uncertainties.
@@
    By compressing high-dimensional channel data into latent representations and generating synthetic training samples, GenAI models reduce computational overhead while simultaneously enhancing resource allocation, beam tracking, and energy efficiency. This dual capability of optimizing performance while minimizing complexity is crucial for real-time operations in resource-constrained LAE scenarios.
@@
    The ability of GenAI models to to model evolving threats and environmental dynamics supports proactive security. By synthesizing attack patterns and channel anomalies, GenAI models can generate adaptive anti-jamming strategies while maintaining reliable beamforming performance, which is essential for mission-critical LAE applications such as emergency communications.
@@ IV-A Motivation and System Design
In LAE scenarios, security risks are rising due to complex terrains and high-density LAP deployments, which often lead to vulnerabilities in communication, with traditional encryption and anti-jamming methods struggling under resource constraints [15]. Additionally, energy efficiency is critical for LAPs due to their limited power supply and high energy consumption. To address these challenges, this case study utilizes UAVs, key components of LAPs, to implement collaborative aerial beamforming for secure remote communications. Specifically, multiple UAVs form a virtual antenna array to enhance the signal directionality and strength towards the remote base station, while optimizing excitation current weights and UAV positions to ensure secure and energy-efficient data transmission. The proposed optimization framework combines the GDM model with DRL to improve both security and energy efficiency.
IV-B Proposed Algorithm
To address the challenges identified in the system design, this study proposes the GDM-enhanced twin delayed deep deterministic policy gradient (GDMTD3) algorithm, which integrates a GDM into the twin delayed deep deterministic policy gradient (TD3) framework. This integration allows the algorithm to optimize UAV swarm actions by modeling the problem as a Markov decision process. Specifically, the state is defined by the positions of the UAVs and the estimated location of the eavesdropper. Based on this, the action consists of selecting excitation current weights and adjusting UAV positions. Furthermore, the reward function combines the secrecy rate and energy consumption, while also imposing penalties for violations such as exceeding speed limits or UAV collisions.
@@ IV-B Proposed Algorithm
IV-C Performance Evaluation
@@ V-B Long-Range High-Power Wireless Power Transfer-Driven Beamforming
Unlike short-range wireless power transfer (WPT) targeting low-power devices and sensors, LAPs require long-range, high-power WPT for reliable and sustainable operation. In this context, it is necessary to realize long-range high-power WPT with beamforming, such as laser and RF-based power beamforming. Moreover, integrating space-based solar power with such advanced WPT solutions can address energy challenges, ultimately supporting the development of 6G systems across various applications in LAE.

### v2

@@ III-B1 Communication Tasks
  • Channel Modeling and Estimation: Compared to conventional communication scenarios, LAE places higher demands on channel modeling and estimation due to its complex and highly dynamic environment. Specifically, factors such as adverse weather conditions, obstruction by buildings or terrain features, interference from surrounding nodes, and the rapid mobility of LAPs significantly impact signal quality. These factors often lead to signal attenuation and pronounced multipath propagation effects, thus posing substantial challenges to accurate channel modeling and estimation. Although conventional beamforming techniques can adapt to certain channel variations, they struggle to effectively address the rapid temporal and spatial fluctuations introduced by these environmental factors. Therefore, there is a clear need for innovative approaches, and GenAI offers promising solutions by leveraging advanced data-driven frameworks to improve channel modeling and estimation.
    For instance, Zhang et al. [11] developed a GAN-based model to improve channel estimation in millimeter-wave communications for low-altitude unmanned aerial vehicle (UAV) networks. Specifically, the model allows each UAV to independently train channel models for beamforming directions, with samples shared through a collaborative framework to build a generalized model, which improves estimation accuracy. Additionally, Rasheed et al. [12] proposed a distributed GAN model based on long short-term networks for low-altitude maritime UAV communications. Specifically, UAVs utilize the model to extract real-time CSI and generate synthetic samples, which improve the accuracy of channel state estimation.
@@ III-B2 Performance Metrics
  • Energy Consumption: In LAE senarios, the frequent beam redirection by LAPs, interference suppression between high-density nodes, and multi-network handovers significantly increase LAP energy consumption compared to conventional communication environments. However, conventional beamforming algorithms rely on pre-set power distribution strategies, which hinder the ability to balance communication quality and energy efficiency in dynamic topology and channel fluctuation scenarios. GenAI can learn the relationship between environmental features and energy consumption online, autonomously generating beam control strategies that optimize energy efficiency and overcome the limitations of traditional static approaches.
@@ III-C Lessons Learned
    GenAI models fundamentally enhance the adaptability, inherent robustness, and real-time responsiveness in beamforming processes, particularly in complex, dynamic LAE environments. Specifically, the capacity to learn underlying data distributions, quantify uncertainty, and efficiently infer from complex inputs enables rapid adaptation to changing channel conditions and resilient data reconstruction even under severe signal degradation. Moreover, the operational agility is vital for maintaining high-quality, continuous communication and precise tracking across diverse data types, including text, images, and video, thus ensuring robust performance despite environmental uncertainties and mobility constraints.
@@
    GenAI significantly improves channel modeling and estimation in LAE environments. Specifically, GenAI models excel at learning intricate channel characteristics even with limited data, thus enabling accurate real-time CSI extraction. In addition, GenAI models also fundamentally redefine resource allocation. In particular, GenAI-enhanced beamforming methods can simultaneously adjust signal directions and spectrum assignments. Consequently, the capability of GenAI ensures efficient resource management and effectively mitigates interference in highly dynamic conditions of LAE.
@@
    For communication security, GenAI models offer significant advancements in LAE. Specifically, GenAI models adaptively optimize anti-jamming strategies against dynamic threats, which are crucial for critical missions [15]. Moreover, GenAI provides solutions for energy consumption. In particular, GenAI-based models learn environmental features online to generate optimized, energy-efficient beam control strategies, thereby maintaining performance while reducing power use.
@@ IV-A Motivation and System Design
In LAE scenarios, security risks are rising due to complex terrains and high-density LAP deployments, often leading to communication vulnerabilities where traditional encryption and anti-jamming methods struggle under resource constraints. Therefore, high secrecy rates are crucial to prevent eavesdropping on sensitive missions in open airspace. Additionally, minimizing energy consumption is vital for long-term operations of LAPs, given their limited onboard batteries and high energy demands. Moreover, a tradeoff exists between secrecy rate and energy consumption. Specifically, boosting security might mean LAPs fly longer or more complex paths, increasing energy use. Conversely, low battery levels could prevent energy intensive security maneuvers. As a result, joint optimization of both is essential for a practical balance.
As the precise synchronization and high-precision localization technologies for distributed aerial platforms have matured, collaborative beamforming has emerged as an effective means of achieving long-distance secure communications. Therefore, this case study utilizes UAVs, key components of LAPs, to implement collaborative aerial beamforming for secure remote communications. Specifically, multiple UAVs form a virtual antenna array to enhance the signal directionality and strength towards the remote base station, while optimizing excitation current weights and UAV positions to ensure secure and energy-efficient data transmission. The proposed optimization framework combines the GDM model with DRL to improve both security and energy efficiency.
IV-B Simulation Setup:
In this section, we detail the computational environment and environmental parameters used to evaluate our proposed framework.
  • Simulation Platform: Our experiments are conducted using a computing setup that included an NVIDIA GeForce RTX 3090 GPU with 24 GB of memory and a 13th Gen Intel(R) Core(TM) i9-13900K 32-core processor with 128 GB of RAM. The operating system on the workstation is Ubuntu 22.04.3 LTS. For our deep learning computations, we use PyTorch 2.6.0, along with the CUDA 12.4.
  • Environmental Details: In this study, we consider a UAV swarm consisting of 4 units, each equipped with a transmit power of 0.1 W. Moreover, the swarm is randomly distributed within a 40 m × 40 m area. To simulate potential security threats, a mobile eavesdropper is introduced, which follows the Gauss-Markov mobility model with an average speed of 5.0 m/s, a correlation coefficient of 0.1, and a variance of 1.0.
IV-C Proposed Algorithm
Considering the challenges posed by the dynamic and uncertain nature of LAE environments, common DRL algorithms often struggle to achieve stable and proper performance. To address these limitations, this study proposes the GDM-enhanced twin delayed deep deterministic policy gradient (GDMTD3) algorithm, which integrates a GDM into the twin delayed deep deterministic policy gradient (TD3) framework to enhance the comprehension of environmental dynamics. This integration allows the algorithm to optimize UAV swarm actions by modeling the problem as a Markov decision process. Specifically, the state is defined by the positions of the UAVs and the estimated location of the eavesdropper. Based on this, the action consists of selecting excitation current weights and adjusting UAV positions. Furthermore, the reward function combines the secrecy rate and energy consumption, while also imposing penalties for violations such as exceeding speed limits or UAV collisions.
@@ IV-C Proposed Algorithm
In addition to the above, the proposed system also possesses significant resilience and broad applicability across diverse environments. In particular, the GenAI-enabled framework integrated into system design adapts to varying device mobility and heterogeneous antenna configurations through dynamic learning capability. Moreover, decoupled offline training ensures compatibility with resource-constrained edge deployments. Furthermore, inherent robustness allows GDMTD3 to dynamically handle abnormal situations by perceiving such occurrences as environmental changes and then adjust policies accordingly. For example, if a UAV in the swarm experiences a fault, the system immediately perceives this as an environmental change. Consequently, GDMTD3 algorithm then dynamically adjusts the communication parameters, such as re-routing data traffic through other functional UAVs or reallocating tasks to more reliable nodes, to maintain robust data links and ensure reliable operation despite the failure of individual UAVs.
IV-D Performance Evaluation
@@ V-B Long-Range High-Power Wireless Power Transfer-Driven Beamforming
Unlike short-range wireless power transfer (WPT) targeting low-power devices and sensors, LAPs require long-range, high-power WPT for reliable and sustainable operation. In this context, it is necessary to realize long-range high-power WPT with beamforming, such as laser and radio frequency (RF)-based power beamforming. Moreover, integrating space-based solar power with such advanced WPT solutions can address energy challenges, ultimately supporting the development of 6G systems across various applications in LAE.